OpenStudy (anonymous):

You join a paperback book club for $6 and then spend $3 per book. Y = 3x + 6. Type how to graph the equation using the y-intercept and slope. Be sure to tell where you place the first point and how to find the location of the second point. really dont understand how

hero (hero):

The equation Y = 3x + 6 is written in the form Y = mx + b where m = slope and b = y-intercept. The y-intercept is 6. This means that the first point you graph will be (0,6). To graph the next point, use the slope m = 3/1.

OpenStudy (anonymous):

Thanks :)